the number of real solution of the equation


|x|^2 - 3|x| + 2 = 0 is


1. 4


2. 1


3. 3


4. 2

Grade:11

1 Answers

Rishabh Jain
13 years ago

for x≥0

x2-3x+2=0 comparing with ax2+bx+c=0,

discriminant>0

=>it has 2 real solutions

now for x≤0

x2+3x+2=0

similarly, it has also 2 solutions

total sol.=4